FTIR AND ULTRAFAST INFRARED SPECTROSCOPY OF THE DICYANAMIDE ANION IN SOLUTION

\underline{KEVIN DAHL}, GERALD M. SANDO and JEFFREY C. OWRUTSKY, Chemistry Division, U.S. Naval Research Laboratory, Washington, DC 20375.

Steady state and time-resolved IR spectroscopy have been used to study the vibrational band positions, vibrational energy relaxation (VER) rates, and reorientation times of the antisymmetric C≡N stretching band of dicyanamide (N(CN)_{2}^{−}) anion near 2100 cm^{−1} in several bulk solvents. The transient signals indicate a complicated relaxation mechanism, consistent with there being a Fermi resonance enhanced combination band in the region. The solvent-induced spectral blue shifts and VER rates both increase for solvents with stronger solute-solvent interactions. The gas-phase vibrational frequency is estimated by extrapolating to zero solvent shift by scaling the solvent shift to those previously observed for azide (N_{3}^{−}). N(CN)_{2}^{−} exhibits a strong correlation between vibrational frequency and VER rate, so that, as with N_{3}^{−}, an estimate of the relative VER rate can be obtained from the static spectra.